OpenStudy (anonymous):

Can anyone teach me how to do the Laplace transform of \(t^2+1\)

OpenStudy (anonymous):

you mean derivation or simply using the rules they supply you after the rules

OpenStudy (anonymous):

after the erivation*

OpenStudy (anonymous):

i guess i am just not familiar with integration from 0 to \(\infty\)

OpenStudy (anonymous):

\[\int _0 ^\infty e^{-st}(t^2+1)dt\]

OpenStudy (anonymous):

\[\int e^{-st} t^2dt+\int e^{-st}\]

OpenStudy (anonymous):

oh ok i get it now i was doing them both together...

OpenStudy (anonymous):

thanks lol -_- im dumb